Solve the problem. Assume that simple interest is being calculated in each case. Round your answer to the nearest cent.

10) Annie's cafe borrows $4700 at 9% for 80 days. Find the total amount that must be repaid after 80 days. (Use a 365 day year.)
Answer #1

P=4700

R=9%

T=80days = 80/365 year

total amount = P + P*R*T/100 = 4700+4700*(80/365)*(9/100)= $4792.71